OpenStudy (anonymous):

An equilateral triangle of side 5 square root3 has its sides increasing in length at the same rate. if the area of the triangle is increasing at 6cm^2/minute, find the rate at which a side is increasing in length.

OpenStudy (anonymous):

If the side has length L cm its area A= ½L²sin60°=(√3/2)L² cm² dA/dt=√3LdL/dt=6 so dL/dt=(6/√3)(1/L)=(2√3)/L So you need to know the value of L and if L=5√3 cm then dL/dt =2√3/(5√3)=2/5 cm/minute.

OpenStudy (anonymous):

thx :) ... is it a problem if u explain the second pls cause i couldn't understand it :) ?